OpenStudy (anonymous):

Use half angle identities to fine the exact value for sin 5pi/12

OpenStudy (jdoe0001):

$$\bf \large { \cfrac{5\pi}{12} \times 2 \implies \cfrac{5\pi}{6}\\ \cfrac{\frac{5\pi}{6}}{2} \implies \cfrac{5\pi}{12}\\ sin\left(\cfrac{\frac{5\pi}{6}}{2}\right) = \pm\sqrt{\frac{1-cos\left(\frac{5\pi}{6}\right)}{2}} } $$